Three pens and three notebooks cost $6.90. Two notebooks and one pen cost $4.10. Model this by useing a system of linear equations
Answer by ikleyn(52781)

.
Three pens and three notebooks cost $6.90. Two notebooks and one pen cost $4.10. Model this by useing a system of linear equations
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

3p + 3n = 690,   (1)
1p + 2n = 410    (2)

You asked to model it. I modelled it for you.


But let me add couple of words to it.

If "Three pens and three notebooks cost $6.90", then 
    one pen    and one   notebook cost one third of it, i.e. 690/3 = 230 cents, right ?

Next, if one pen and one notebook  cost 230 cents    and
         one pen and two notebooks cost 410 cents,   THEN WHAT ?

But of course, then one notebook costs 410 - 230 = 180 cents = $1.80.


So, one half of the problem is just solved.


It is how the Elimination Method works.
I explained it to you by using words, without equations.

Happy learning !